# Date Fragments: A Hidden Bottleneck of Tokenisation for Temporal Reasoning

### v2

@@ 8 Conclusion
While our work demonstrates the impact of date tokenisation on LLMs for temporal reasoning, there are several limitations. First, DateAugBench focuses on a finite set of canonical date serialisations and does not capture the full diversity of natural-language expressions (e.g., “the first Monday of May 2025”) or noisy real-world inputs like OCR outputs. Second, our experiments evaluate a representative but limited pool of tokenisers and model checkpoints (up to 14B parameters); therefore, the generalizability of date fragmentation ratio and our probing and causal analyses to very large models with 15B+ parameters remains unknown. Third, while the fragmentation ratio measures front-end segmentation fidelity, it does not account for deeper world-knowledge factors such as leap-year rules, timezone conversions, and culturally grounded calendar systems, all of which may influence temporal interpretation; further, the fragmentation ratio metric, though straightforward and interpretable, is not rigorously evaluated. Lastly, the core idea of our causal framework is inspired by Lindsey et al. (2025); however, our extension to temporal reasoning is not evaluated. Future work should extend to more diverse date expressions, broader model and tokeniser families, equipping tokenisers with external calendar-wise knowledge to improve further robust temporal reasoning, and conducting rigorous evaluation of the fragmentation ratio metric and the causal framework.
@@ Acknowledgements
We gratefully thank Madiha Kazi, Cristina Mahanta, and MingZe Tang for their support of conducting human evaluation for LLMs-as-judge. We also thank Ahmad Isa Muhammad for participating in early discussions.
@@ A.2.1 Next Token Prediction
We treat each token in the prompt as a candidate “concept” to follow. After the model processes the input, it produces a hidden vector $h_{\ell,p}$ per token at position $p$ and layer $\ell$ To see how likely a concept $c$ (e.g., a date fragment and model answer) follows each input token, we project $h_{\ell,p}$ through $W_{U}$ to yield the “probability” distribution of vocabulary tokens, and denote $s^{c}_{\ell,p}$ as the “probability” of the concept being the next token.
@@ A.2.3 Path scoring
  • Ordering: we give points if the concepts appear in roughly left-to-right order in the prompt, and secondarily in increasing layer order:
@@
Model DD-MM-YYYY DD/MM/YYYY YYYY/MM/DD DDMMYYYY MMDDYYYY YYYYMMDD Avg.
@@
Figure 9: Reasoning path of the “03121325 is a valid date” prompt, where year $=1325$.
@@
LLM-as-Judge Evaluation Prompt Your task: Evaluate one prediction at a time. You receive: Question – the task prompt shown to the model Gold targetall answers that are considered correct Predicted answer – the model’s response Return one letter only: A CORRECT prediction fully matches one gold variant B INCORRECT prediction contradicts or misses required info C NOT_ATTEMPTED prediction refuses, guesses, or answers irrelevantly General rules: 1. Match semantics, ignore capitalisation, punctuation, order. 2. If any statement contradicts the gold target, grade B. 3. Hedging ("I think…") is fine if the correct info is present and no incorrect info is added. 4. Partial answers are B. Typos that preserve meaning are allowed. DateAugBench specifics: Date format ambiguity: gold lists every valid interpretation; accept any. Date arithmetic: prediction must match day, month, year of a listed variant, any textual format allowed. Format-switch questions: answer with any synonym of Yes/True or No/False. Numeric answers – must match the gold number to the last shown significant digit. Output format

### v3

@@ 9 Conclusion
While our work demonstrates the impact of date tokenisation on LLMs for temporal reasoning, there are several limitations. First, DateAugBench focuses on a finite set of canonical date serialisations and does not capture the full diversity of natural-language expressions (e.g., “the first Monday of May 2025”) or noisy real-world inputs like OCR outputs. Second, our experiments evaluate a representative but limited pool of tokenisers and model checkpoints (up to 14B parameters); therefore, the generalizability of date fragmentation ratio and our probing and causal analyses to very large models with 15B+ parameters remains unknown. Third, while the fragmentation ratio measures front-end segmentation fidelity, it does not account for deeper world-knowledge factors such as leap-year rules, timezone conversions, and culturally grounded calendar systems, all of which may influence temporal interpretation; further, the fragmentation ratio metric, though straightforward and interpretable, is not rigorously evaluated. Our work and the DateAugBench benchmark deliberately focus on a specific, foundational challenge: the tokenisation of explicit, multi-digit date strings in Anglo-centric Gregorian formats. This narrow scope allows us to isolate the impact of subword fragmentation on core temporal reasoning. However, this focus means we do not address the full complexity of temporal understanding, such as dates expressed in natural language (e.g., "the first Monday of May 2025"), dates with missing components, non-Gregorian calendars (e.g., Hijri, Hebrew), or dates represented with non-Latin numeral systems. We consider the extension to these diverse and important cases as critical future work that can build upon our foundational analysis of fragmentation. Lastly, the core idea of our causal framework is inspired by Lindsey et al. (2025); however, our extension to temporal reasoning is not evaluated. Future work should extend to more diverse date expressions, broader model and tokeniser families, equipping tokenisers with external calendar-wise knowledge to improve further robust temporal reasoning, and conducting rigorous evaluation of the fragmentation ratio metric and the causal framework.
@@ Acknowledgements
We thank the anonymous reviewers for their thoughtful comments that greatly improved the work. We gratefully thank Madiha Kazi, Cristina Mahanta, and MingZe Tang for their support of conducting human evaluation for LLMs-as-judge. We also thank Ahmad Isa Muhammad for participating in early discussions.
@@ A.2.1 Next Token Prediction
We treat each token in the prompt as a candidate “concept” to follow. After the model processes the input, it produces a hidden vector $h_{\ell,p}$ per token at position $p$ and layer $\ell$ . To see how likely a concept $c$ (e.g., a date fragment and model answer) follows each input token, we project $h_{\ell,p}$ through $W_{U}$ to yield the “probability” distribution of vocabulary tokens, and denote $s^{c}_{\ell,p}$ as the “probability” of the concept being the next token.
@@ A.2.3 Path scoring
  • Ordering: we give points if the concepts appear in roughly left‐to‐right order in the prompt, and secondarily in increasing layer order:
@@
A.3 Detailed Validation of the Date Fragmentation Ratio
This appendix provides a detailed account of the two-part validation process for our custom date fragmentation ratio (F), demonstrating its alignment with human intuition and its empirical soundness.
A.3.1 Human Evaluation of Fragmentation Severity
This study was designed to confirm that our F metric captures what humans perceive as semantic disruption in tokenized dates more effectively than general-purpose text similarity metrics.
Methodology.
We recruited five computer science graduate students, who were familiar with NLP but blind to our hypotheses, to serve as annotators. We created a stimulus set of 100 tokenised date strings, stratified to represent a wide range of models, date formats, and fragmentation levels from our experiments. For each item, annotators were shown the original date and the list of sub-tokens, and asked to rate the “fragmentation severity” on a 5-point Likert scale, according to the following rubric with examples:
  • 1 (No Fragmentation): Tokens perfectly preserve the semantic components. Example: ‘10-27-1606‘ $\rightarrow$ ‘[’10’, ’-’, ’27’, ’-’, ’1606’]‘
  • 2 (Minor Fragmentation): Mostly preserved, with minor, non-ideal splits. Example: ‘1606‘ $\rightarrow$ ‘[’16’, ’06’]‘
  • 3 (Moderate Fragmentation): Core components are broken, making the structure harder to discern. Delimiters might be lost or numbers oddly grouped. Example: ‘10271606‘ $\rightarrow$ ‘[’102’, ’716’, ’06’]‘
  • 4 (High Fragmentation): Date split into many small pieces (e.g., single digits), though the original characters are easily reassembled. Example: ‘1606‘ $\rightarrow$ ‘[’1’, ’6’, ’0’, ’6’]‘
  • 5 (Severe Fragmentation): Tokenization completely obscures the date’s structure, often by adding non-numeric tokens or creating highly unintuitive groupings. Example: ‘1606‘ $\rightarrow$ ‘[’_’, ’1’, ’6’, ’0’, ’6’]‘
The human judgments were highly reliable, with a Krippendorff’s Alpha for inter-annotator agreement of $\alpha=0.81$.
Results.
We computed the Spearman’s rank correlation coefficient ($\rho$) between the average human rating for each item and the scores from our F metric, BLEU, and character-level Edit Distance. As shown in Table 6, our F metric demonstrated a strong correlation with human ratings, far exceeding the general-purpose metrics. In Table 5, we present examples from the human validation study.
Model Tokenised output Frag-ratio Avg. Human Severity Rating (1-5)
Baseline 10 27 1606 0.00 1.0
OLMo 10 27 16 06 0.34 2.0
Llama 3 102 716 06 0.40 3.4
GPT-4o 102 716 06 0.40 3.4
Cohere 1 0 2 7 1 6 0 6 0.55 4.6
Phi 3.5 _ 1 0 2 7 1 6 0 6 0.60 5.0
Llama 2 _ 1 0 2 7 1 6 0 6 0.60 5.0
Table 5: An illustrative example showing the strong correlation between the calculated fragmentation ratio (Frag-ratio) and the average human-perceived severity rating for the tokenisation of the MMDDYYYY string "10271606". Higher scores in both metrics indicate greater fragmentation.
Metric Correlation with Human Ratings ($\rho$)
Date Fragmentation Ratio (F) 0.84
BLEU Score 0.52
Character-Level Edit Distance 0.21
Table 6: Spearman Correlation ($\rho$) of Metrics with Human Judgments of Fragmentation Severity.
This result confirms that our specialised metric is necessary and effective, as it successfully quantifies the semantic disruption that humans perceive but that generic text metrics fail to capture.
A.3.2 Data-Driven Validation of Metric Coefficients
This analysis provides an empirical grounding for the weights used in our F metric’s formula. By learning the weights from data, we can validate that our initial, intuitive design aligns with a more formal, data-driven approach.
Formal Problem Formulation.
To directly tune our metric to align with human perception, we frame the task as a linear regression problem where the goal is to predict the average human severity rating. This setup is more straightforward for validating the metric’s components against human judgment.
  • The target variable is the Average Human Severity Rating, a continuous score from 1 to 5, as described in Appendix A.3.1.
  • The feature vector $\mathbf{x}\in\mathbb{R}^{4}$ consists of the four fragmentation components: $\mathbf{x}=[1_{\text{split}},1_{\text{delimiter}},(N-N_{b}),\theta]$.
  • We aim to learn a weight vector $\mathbf{w}$ such that:
    $\text{Avg. Human Severity Rating}\approx\mathbf{w}^{T}\mathbf{x}+\text{intercept}$.
We used a non-negative linear regression model, as each fragmentation component is hypothesised to increase, not decrease, the perceived severity. Features were standardised before training to ensure the learned coefficients were comparable.
Results and Confirmation.
After fitting the model to our human evaluation data, we obtained a set of empirically derived coefficients. We normalised these weights to sum to 1 to compare them with the relative importance implied by our original formula. As Table 7 shows, the weights learned by predicting human ratings are remarkably similar to the normalised version of our original, intuitively set weights.
Fragmentation Component Original Intuitive Weight Empirically Learned Weight
(Normalized) (from Human Ratings)
$1_{\text{split}}$ (Component Split) $0.10/0.55\approx\textbf{0.1818}$ 0.2015
$1_{\text{delimiter}}$ (Delimiter Loss) $0.10/0.55\approx\textbf{0.1818}$ 0.1932
$N-N_{b}$ (Token Difference) $0.05/0.55\approx\textbf{0.0909}$ 0.1053
$\theta$ (Distributional Divergence) $0.30/0.55\approx\textbf{0.5455}$ 0.5000
Table 7: Comparison of Original (Normalised) and Empirically Learned Weights for the F Metric, using human ratings as the target variable.
This result provides strong empirical validation of our F metric’s design from an alternative perspective. It demonstrates that our initial weights, chosen based on semantic principles, accurately reflect not only the impact on model performance but also the factors that drive human perception of the severity of fragmentation. The relative importance of the components remains consistent: distributional divergence ($\theta$) is the most significant factor, followed by major structural breaks (splits and delimiter loss), and finally by token count inflation. This confirms the robustness and validity of our metric’s formulation.
@@ A.3.2 Data-Driven Validation of Metric Coefficients
Model $D$-$M$-YYYY DD/MM/YYYY YYYY/MM/$D$ DDMMYYYY MMDDYYYY YYYYMMDD Avg.
@@
Figure 9: Reasoning path of the “03121325 is a valid date” prompt, where year =1325=1325.
@@
LLM‑as‑Judge Evaluation Prompt Your task: Evaluate one prediction at a time. You receive: Question – the task prompt shown to the model Gold targetall answers that are considered correct Predicted answer – the model’s response Return one letter only: A CORRECT prediction fully matches one gold variant B INCORRECT prediction contradicts or misses required info C NOT_ATTEMPTED prediction refuses, guesses, or answers irrelevantly General rules: 1. Match semantics, ignore capitalisation, punctuation, order. 2. If any statement contradicts the gold target, grade B. 3. Hedging ("I think…") is fine if the correct info is present and no incorrect info is added. 4. Partial answers are B. Typos that preserve meaning are allowed. DateAugBench specifics: Date format ambiguity: gold lists every valid interpretation; accept any. Date arithmetic: prediction must match day, month, year of a listed variant, any textual format allowed. Format‑switch questions: answer with any synonym of Yes/True or No/False. Numeric answers – must match the gold number to the last shown significant digit. Output format